David and Nicole have a new grandson. How much money should they invest now so that he will have $68000 for his college education in 18 years? The money is invested at 3.8% compounded quarterly.
Answer by mananth(16946)

David and Nicole have a new grandson. How much money should they invest now so that he will have $68000 for his college education in 18 years? The money is invested at 3.8% compounded quarterly.
Amount = $68,000
time = 18 years
rate = 3.8% = 0.038
calculated 4 times a year
Principal P = x
Amount= 68000
years=n 18
compounded 4 times a year
Rate = 3.80 0.038 ... 0.038/4 =0.950
Amount = P*((n+r)/n)^n
1.07
68000 = x *( 1 + r/n )^ n


68000 = P ( 1 + 0.038 / 4 ) ^ 4 * 18

68000 = P( 1.00950 )^ 72

68000 /( 1.00950 )^ 72 = P

P= $34423.39